OpenStudy (lordhelix8th):

could someone help me with 8 questions

OpenStudy (lordhelix8th):

A cylindrical container holding sugar has a height of 17 inches and a diameter of 4 inches. Which of the following is the closest to the volume of the sugar container? A. 214 cubic inches B. 854 cubic inches C. 107 cubic inches D. 907 cubic inches

OpenStudy (anonymous):

i can try

OpenStudy (lordhelix8th):

@F_Jayyy @poopsiedoodle

OpenStudy (lordhelix8th):

i will fan and medal

OpenStudy (f_jayyy):

i dont know sorry :(

OpenStudy (f_jayyy):

im a little rusty on measurements

OpenStudy (lordhelix8th):

@jabez177

OpenStudy (anonymous):

again i could help!

OpenStudy (lordhelix8th):

ok

OpenStudy (anonymous):

whats your questions

OpenStudy (lordhelix8th):

ill post 1 at a time

OpenStudy (lordhelix8th):

but could u solve the 1 up top?

OpenStudy (anonymous):

Okeh

OpenStudy (lordhelix8th):

i think its c

OpenStudy (anonymous):

yeh

OpenStudy (lordhelix8th):

thats my 2nd question

OpenStudy (lordhelix8th):

@TheSmartOne

OpenStudy (lordhelix8th):

anyone?

OpenStudy (lordhelix8th):

@texaschic101

OpenStudy (lordhelix8th):

@LegendarySadist @Astrophysics

OpenStudy (lordhelix8th):

@pooja195

OpenStudy (4everaddicted2anime):

the answer to the second question is C

OpenStudy (lordhelix8th):

how

OpenStudy (4everaddicted2anime):

\[V=\pi * r ^{2} \frac{ h }{ 3 }\] So when you plug in the given numbers you get \[V= \pi *4^{2} \frac{ 3 }{ 3 }\] Then you just simplify the fraction and sqaure 4 to get the answer \[V=16\pi\]

OpenStudy (lordhelix8th):

thanks the next question is above you

OpenStudy (lordhelix8th):

@bibby

OpenStudy (4everaddicted2anime):

I got B for this one

OpenStudy (lordhelix8th):

so b for the cubic unit 1?

OpenStudy (bibby):

V≈1385.44

OpenStudy (lordhelix8th):

ok hold on

OpenStudy (lordhelix8th):

posting next question

OpenStudy (lordhelix8th):

The table shows the results of a random survey of students in grade 7 and grade 8. Every student surveyed gave a response. Each student was asked if he or she exercised less than 5 hours last week or 5 or more hours last week. Less than 5 hours 5 or More Hours Grade 7 Students 49 63 Grade 8 Students 58 51 Which statement is true? A. More grade 8 students were surveyed than grade 7 students B. Less than 50% of the grade 8 students surveyed exercised 5 or more hours last week C. A total of 107 grade 7 students were surveyed D. More than 50% of the students surveyed exercised less than 5 hours last week

OpenStudy (lordhelix8th):

this is a better view of the question

OpenStudy (lordhelix8th):

bibby where u go

OpenStudy (lordhelix8th):

@bibby

OpenStudy (lordhelix8th):

@Ghostgate

OpenStudy (lordhelix8th):

plz help

OpenStudy (lordhelix8th):

@ParthKohli

OpenStudy (ghostgate):

Problem 1(On top): Is B. 854 (Est.). Why? Because to find the volume of a cylinder object, the formula is: V = pi*r^2*h (V being Volume, r being radius, and h being height.) That comes out at 854.08 when calculating for the answer. Problem 2: Is C. 16 * pi Why? Because to find the volume of a cone, the formula is: V = pi*r^2*h/3 (V being Volume, r being radius, and h being height.) When calculated you get: 16 * pi Problem 3: Is B. 1,384.74 cubic units. Back to the cylinder formula: V = pi*r^2*h (V being Volume, r being radius, and h being height.) This calculates out to: 1,384.74 Sorry for not having full descriptions for checking over the problems everyone else went through with you. Alright, now to go over your latest problem.

OpenStudy (ghostgate):

The table shows the results of a random survey of students in grade 7 and grade 8. Every student surveyed gave a response. Each student was asked if he or she exercised less than 5 hours last week or 5 or more hours last week. Less than 5 hours 5 or More Hours Grade 7 Students 49 63 Grade 8 Students 58 51 Which statement is true? A. More grade 8 students were surveyed than grade 7 students. B. Less than 50% of the grade 8 students surveyed exercised 5 or more hours last week. C. A total of 107 grade 7 students were surveyed. D. More than 50% of the students surveyed exercised less than 5 hours last week. Alright, we need to try and prove whether or not these statements are true: So for A. simply add the students together: 49 + 63 = x 58 + 51 = y For B. knowing the total amount of students from when you were proving if A. was true or not, we plug that in: x/2 = ? - Then just simply view to see if 51 is above or below it to verify. For C. Add 49 and 63 together and see if that reaches 107, if not then C. is false. For D. Add all of the students together. 49 + 63 + 58 + 51 = x x/2 = y 49 + 58 = a - Is a higher than 'y' or lower? If you need help, ask and I'll go over the answers.

OpenStudy (lordhelix8th):

i only got 4 more questions

OpenStudy (lordhelix8th):

and thank you

OpenStudy (ghostgate):

Okay, for this problem we need to figure out which statement is true yet again. We will say that the left triangle has points ABC which have coordinates: A(0,2), B(4,7), C(4,3) Now the right triangle has points DEF which have coordinates: D(-2,-4), E(3,0), F(2,-4) So... now to figure out which one of the statements is true or not. So take the coordinates and translate across, I can solve this, but I don't know how to explain it exactly. Reflecting over the y-axis means that the x-axis usually becomes the inverse. Whereas reflecting over the x-axis means the the y-axis usually is inverted. By inverted/inverse I mean positive to negative or negative to positive. Does that help? Also you're welcome.

OpenStudy (ghostgate):

I would guess C., but I could very much be wrong. Sorry I can't be more help.

OpenStudy (lordhelix8th):

its ok

OpenStudy (lordhelix8th):

^thats a question

OpenStudy (ghostgate):

We definitely know that A. is false because no accidents occurred with 6 hours of rain.

OpenStudy (ghostgate):

I think the answer is C. considering that it seems the most logical or truthful, D. is false, and B. can be presumed false.

OpenStudy (lordhelix8th):

i was think c too

OpenStudy (ghostgate):

Alright, 2 more questions left, correct?

OpenStudy (lordhelix8th):

yes

OpenStudy (ghostgate):

Okay this seems more solvable, here's the chart and question: The owner of a food truck collected data on whether customers purchased tacos or burrito and whether they chose chicken or beef for the meat. The two-way table below shows the data collected. Taco Burrito Total Chicken 13 25 38 Beef 37 25 62 Total 50 50 100 Which correlation does the two-way table suggest? A. Customers who purchase a burrito tend to choose beef. B. Customers who purchase tacos tend to choose chicken. C. Customers who purchase a burrito tend to choose chicken. --->D. None of the correlations listed are correct. The answer is D. because as you can see the other 3 statements are false. A and C are false because when people buy a burrito it is 25 for Chicken and 25 for Beef, neither are higher than the other. B. is false because as far as Tacos go they tend to choose beef, not chicken. So that means D. is the correct answer.

OpenStudy (lordhelix8th):

this is the last one and thank you for everything

OpenStudy (ghostgate):

I think the last one is A, but I may be wrong. Have a great day! ;) {---Ghostgate---}